OpenStudy (anonymous):

The width of a rectangle is 6 feet, and the diagonal is 10 feet. What is the area of the rectangle?

OpenStudy (wolf1728):

length² = diag² - 6² lebgth² = 100 - 36 length² 64 length = 8 area = 8 * 6 = 48